How fo you calculate MARCS mid-quarter depreciation for a 500 000 asset put into service in November. Im confused how you would apply the staight line method with mid quarter
Answer #1

Straight line method with mid quarter -

Value of Asset is 500,000

For example - Useful life - 5 years

Service in date - November

Salvage value zero

So, yearly depreciation is 5,00,000/5 = 100,000 = 20%

Being its mid quarter November rate of depreciation is 20/12*1.5 = 2.5%

in other words Per quarter its 20/4 = 5% and its mid quarter So, 5/2 = 2.5 %